OpenStudy (reneeshumpert):

which of the following is equivalent to (5x+2)^3? A. (25x^2+20x+4)(5x+2) B. (16x^2-8x+1)(4x-1) C. (125x^3-120x^2-60x-8 D. (25x^2-20x-4)(5x+2)

OpenStudy (anonymous):

a or d

OpenStudy (reneeshumpert):

I really wanted somewhat of a walkthrough if thats cool

OpenStudy (faiqraees):

(5x+2)^3 = (5x+2)(5x+2)(5x+2) Multiply the first two expressions and check whether any of the given options is correct

OpenStudy (reneeshumpert):

Give me one second

OpenStudy (reneeshumpert):

None of them match up I got 125x^3+150x^2+60x+8

OpenStudy (faiqraees):

(5x+2)(5x+2)(5x+2) = (5x+2)(5x+2) * (5x+2) Just multiply (5x+2)(5x+2) and put it as "answer"(5x+2)

OpenStudy (reneeshumpert):

So it would be A. ? could you just explain how thats equivalent for me so I can have an understanding. I have a lot of questions like this one and I would like to be able to do them on my own next time.

OpenStudy (faiqraees):

(25x^2+20x+4)(5x+2) = 125x^3+150x^2+60x+8 (5x+2)^3 = 125x^3+150x^2+60x+8 And see it like this (5x+2)^3 says multiply (5x+2) by itself three times (25x^2+20x+4)(5x+2) says multiply (5x+2) by (25x^2+20x+4) which is actually (5x+2)^2

OpenStudy (reneeshumpert):

Thank you a lot, that was a lot of help I gave you a medal hopefully others who come across the same problem will have an understanding as well.!